OpenStudy (anonymous):

What function has a y-intercept of 1/2?

OpenStudy (anonymous):

Slope intercept form y = mx + b where m is the slope and b is the y-intercept Y-intercept: the y-value when x=0

OpenStudy (anonymous):

out of \[f(x)=\frac{ 2 }{ (2x-1)(x+1) }\] \[f(x)=\frac{ 2 }{ 2x ^{2}+5x-3 }\] \[f(x)=-\frac{ 4 }{ x ^{2}-7x-8 }\]

OpenStudy (raden):

plug x=0, u will get the y-intercept value

OpenStudy (anonymous):

thanks you guys!